OpenStudy (anonymous):

Input in standard form the equation of the given line. The line through (0, -3) and (3, 0) Pleas show steps

OpenStudy (anonymous):

Can use intercept-intercept formula: \(\large \frac{x}{a}+\frac{y}{b}=1\), where a = x-intercept, and b=y-intercept. Then rearrange that to look like \(\large Ax+By=C\).

OpenStudy (anonymous):

Find the Slope First: \[m=\frac{ y2-y1 }{ x2-x1 }\] then use Point Slope Form...\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

Ok that makes sense I got the slope of 1 and the equation 0-3=1x(0-3) is this correct?

OpenStudy (anonymous):

No, you need to include x and y in the equation.

OpenStudy (anonymous):

Start with \(\large \frac{x}{3}+\frac{y}{-3}=1\) , intercept-intercept form. Multiplying both sides by the greatest common denominator, 3, to clear the fractions yields \(\large x-y=3\).

OpenStudy (anonymous):

3x+3y=b?

OpenStudy (anonymous):

No, x-y=3.

OpenStudy (anonymous):

oh so is that it?

OpenStudy (anonymous):

If you have both intercepts of a line, using intercept-intercept form gets you to standard form in one step, yes. You can also do what Yahoo! suggests, as that is more general and works for any given points, but requires many more steps. I recommend doing it both ways to confirm that they are equivalent (and for the extra practice).

OpenStudy (anonymous):

Thanks